# VidCRAFT3: Camera, Object, and Lighting Control for Image-to-Video Generation

### v3

@@ Abstract
Recent image-to-video generation methods have demonstrated success in enabling control over one or two visual elements, such as camera motion or object motion. However, these methods are unable to offer control over multiple visual elements due to limitations in data and network efficacy. In this paper, we introduce VidCRAFT3, a novel framework for precise image-to-video generation that enables control over camera motion, object motion, and lighting direction simultaneously. VidCRAFT3 integrates three core components: Image2Cloud generates 3D point cloud from a reference image; ObjMotionNet encodes sparse object trajectories using multi-scale optical flow features; and Spatial Triple-Attention Transformer incorporates lighting direction embeddings via parallel cross-attention modules. Additionally, we introduce the VideoLightingDirection dataset, providing synthetic yet realistic video clips with accurate per-frame lighting direction annotations, effectively mitigating the lack of annotated real-world datasets. We further adopt a three-stage training strategy, ensuring robust learning even without joint multi-element annotations. Extensive experiments show that VidCRAFT3 produces high-quality video content, outperforming state-of-the-art methods in control granularity and visual coherence. Code and data will be publicly available.
@@ 1 Introduction
Image-to-video (I2V) generation is a powerful technique that brings still images to life, enabling widespread applications in the fields of content creation, advertising, and art animation. Typically, I2V generation aims to animate a reference image according to user-defined control signals, such as text, object motion, and camera motion, while maintaining the quality and natural dynamics of the generated video. Driven by advances in diffusion-based generative models and extensive web-scale data [2], recent methods have significantly enhanced the ability to generate temporally coherent and visually compelling videos from limited inputs such as a single image or sparse annotations [15, 6, 60, 14, 57, 17, 28, 13]. However, achieving precise and simultaneous control over multiple video attributes, such as camera motion, object motion, and lighting direction, remains a critical yet challenging objective.
@@
Existing approaches typically address these control attributes independently or in a partially integrated manner. Methods focusing exclusively on camera control, such as CameraCtrl [17] and ViewCrafter, do not adequately handle detailed object motion or dynamic lighting variations. Similarly, object-motion-centric frameworks, including DragAnything [58], lack integrated capabilities for precise camera and lighting control. For lighting control, techniques like DILightNet [66] and NeuLighting [27] provide lighting control but often lack generalizability due to reliance on specific categories (e.g., human faces) and object-specific annotations. Other methods [27, 68, 22, 68, 22] also primarily target narrow domains, such as human faces, further restricting their applicability to open-domain scenarios.
@@
Simultaneous and fine-grained control of camera motion, object motion, and lighting direction introduces multiple significant challenges: (1) Accurate camera motion control from single-image inputs requires robust extraction of sufficient 3D geometric information. (2) Realistic and detailed object motion control demands effective representation and refinement of sparse motion trajectories without compromising visual fidelity. (3) Dynamic lighting control necessitates integrating illumination adjustments coherently with both camera and object motion to ensure temporal consistency and visual realism.
@@
To overcome these challenges, we propose VidCRAFT3, which integrates three core components. First, the Image2Cloud module leverages DUSt3R [54] to reconstruct a 3D point cloud from a single reference image, enabling precise camera motion control by rendering the point cloud along user-defined camera motion trajectory. Second, ObjMotionNet encodes sparse object motion trajectories by extracting multi-scale motion features from Gaussian-smoothed optical flow maps to guide realistic object motion. Third, Spatial Triple-Attention Transformer integrates lighting direction control by integrating lighting embedding with image-text embeddings through parallel cross-attention layers, ensuring consistent illumination effects. The capabilities of these modules are shown in Fig. 1, showcasing VidCRAFT3’s control over camera motion, object motion, and lighting direction. To address data scarcity, we introduce the VideoLightingDirection (VLD) dataset, providing synthetic yet highly realistic video sequences with accurate per-frame lighting annotations. Additionally, we develop a three-stage training strategy that enables effective learning without joint multi-element annotations.
@@
The main contributions of this paper are: (1) VidCRAFT3 is the first I2V framework to achieve simultaneous control over camera motion, object motion, and lighting direction through a disentangled architecture combining Image2Cloud, ObjMotionNet, and Spatial Triple-Attention Transformer. (2) VLD dataset provides synthetic videos accompanied by accurate per-frame lighting direction annotations, effectively addressing the scarcity of annotated real-world datasets. (3) A three-stage training strategy enables robust multi-element control without requiring real-world videos annotated with camera, object, and lighting signals. (4) Extensive experiments demonstrate that VidCRAFT3 achieves state-of-the-art performance surpassing existing methods in terms of control precision, visual quality, and generalization capabilities.
2 Related Work
2.1 Image-to-video Generation
@@
Image-to-video (I2V) generation [15, 6, 7, 60, 14, 16] aims to animate static images into dynamic videos while preserving visual content and introducing realistic motion. Recent advances in diffusion models [48, 20] have revolutionized video generation by extending pre-trained Text-to-Image (T2I) models like AnimateDiff [15] to incorporate temporal dimensions for motion generation. These methods integrate the input image as a condition, either through CLIP-based [42] image embeddings or by concatenating the image with noisy latent. For example, VideoCrafter1 [6], DynamiCrafter [60], and I2V-Adapter [14] use dual cross-attention layers to fuse image embeddings with noisy frames, ensuring spatial-aligned guidance. Similarly, Stable Video Diffusion (SVD) [4] replaces text embeddings with CLIP image embeddings, maintaining semantic consistency in an image-only manner. Another line of work, exemplified by SEINE [8], DynamiCrafter [60] and PixelDance [67], expands the input channels of diffusion models to concatenate the static image with noisy latents, effectively injecting image information into the model. However, these methods preserve input image fidelity while generating dynamic videos but often struggle with fine-grained details due to reliance on global conditions.
2.2 Motion-controlled Video Generation
Motion-controlled video generation [57, 17, 64] focuses on creating high-fidelity videos with user-defined motion dynamics. Existing methods can be categorized into camera motion control, object motion control, and joint motion control. For camera motion control [17, 1, 26, 21, 56], CamI2V [72], CameraCtrl [17], and CamCo [61] encode camera parameters into Plücker embeddings for more discriminative motion representation. ViewCrafter [65] and I2VControl-Camera [11] use 3D representations of reference images to render point clouds and generate partial frames. For object motion control [32, 46, 55, 23, 30, 71, 74, 62, 37, 19, 12, 34, 59], existing methods employ diverse strategies to guide object movements in video synthesis. DragNUWA [64], Image Conductor [28], DragAnything [58], and MotionBridge [50] utilize sparse optical flow to control object motion, while MOFA-Video [36] and Motion-I2V [45] predict dense motion fields to control object motion. Alternatively, Boximator [52] and Direct-a-Video [63] rely on bounding box trajectories to guide object motion. LeviTor [51] combines depth and K-means clustering for precise 3D trajectory control. Current research on Joint Motion Control [63, 57, 13, 9], aiming to simultaneously control camera and object motions, remains limited. MotionCtrl [57] designs specialized modules for precise control, while Motion Prompting [13] uses motion tracks for flexible guidance. Perception-as-Control [9] further enhances control with 3D-aware motion representations.
@@ 2.2 Motion-controlled Video Generation
Unlike existing methods, we propose VidCRAFT3, the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction. By combining 3D point cloud rendering, trajectory learning, and Spatial Triple-Attention Transformer, our approach effectively decouples these elements, ensuring temporal consistency and enhanced realism in complex scenes.
2.3 Lighting-controlled Visual Generation
@@
Lighting-controlled visual generation aims to manipulate illumination while preserving scene geometry and materials. Previous methods primarily focus on portrait lighting [49, 73, 43, 38, 35, 25, 44, 47, 18, 40, 70], laying the foundation for effective and accurate illumination modeling. Recent advances in diffusion models have significantly improved the quality and flexibility of lighting control. Methods like DiLightNet [66] and GenLit [3] achieve fine-grained and realistic relighting through radiance hints and SVD, respectively. Facial relighting methods, including DifFRelight [18] and DiFaReli [40], produce high-quality portrait images. Frameworks like NeuLighting [27] focus on outdoor scenes using unconstrained photo collections, while GSR [39] combines diffusion models with neural radiance fields for realistic 3D-aware relighting. IC-Light [69] proposes imposing consistent light transport during training. Extending lighting control to video [70] introduces challenges such as temporal consistency and dynamic lighting effects. Recent techniques leverage 3D-aware generative models for temporally consistent relighting, as seen in EdgeRelight360 [29] and ReliTalk [41]. Neural rendering approaches [5, 68] use datasets like dynamic OLAT for high-quality portrait video relighting, while reflectance field-based methods [22] infer lighting from exemplars. Despite these advancements, most prior works focus on portraits and HDR lighting conditions. In contrast, VidCRAFT3 targets general scenes, enabling interactive adjustment of lighting directions.
3 Method
3.1 Overview
@@
Figure 2: Architecture of VidCRAFT3 for controlled image-to-video generation. The model builds on Video Diffusion Model (VDM) and consists of three main components: the Image2Cloud reconstructs 3D point cloud from a single reference image and generates point cloud renderings along user-defined camera motion trajectory; the ObjMotionNet injects object dynamics into the UNet by encoding sparse trajectories into multi-scale motion features; the Spatial Triple-Attention Transformer integrates image, text, and lighting information via parallel cross-attention modules. The model enables I2V generation conditioned on arbitrary combination of camera motion, object motion, and lighting direction.
We present VidCRAFT3, the first framework for high-fidelity Image-to-Video (I2V) generation with disentangled control over camera motion, object motion, and lighting direction. Existing methods struggle to simultaneously manage these elements due to entangled representations and insufficient annotated training data. To address this challenge, as shown in Fig. 2, our framework integrates three core components: (1) Image2Cloud, which reconstructs a 3D point cloud from a single input image and generates point cloud renderings along user-defined camera motion trajectory for precise camera control; (2) ObjMotionNet, which encodes sparse object trajectories into multi-scale motion features for realistic object dynamics; and (3) the Spatial Triple-Attention Transformer, which integrates lighting direction, image, and text features via parallel cross-attention layers. To overcome the scarcity of real-world videos with multi-element annotations, we construct three specialized datasets and adopt a three-stage training strategy to progressively optimize the model.
Our model is built upon the open-source I2V model DynamiCrafter, which consists of a VAE encoder $\mathcal{E}$ and decoder $\mathcal{D}$ for image compression, a UNet with spatial and temporal layers, and CLIP text and image encoders. We employ a dual-stream injection approach where the reference image is encoded by CLIP image encoder and injected into UNet through image cross-attention. Simultaneously, the first frame of the point cloud rendering is replaced with the reference image, which is then encoded by the VAE encoder and concatenated with noise before being fed into the UNet. This design ensures seamless integration of spatial and temporal information for high-quality video generation.
3.2 Model Architecture

### v4

@@ Abstract
Controllable image-to-video (I2V) generation transforms a reference image into a coherent video guided by user-specified control signals. In content creation workflows, precise and simultaneous control over camera motion, object motion, and lighting direction enhances both accuracy and flexibility. However, existing approaches typically treat these control signals separately, largely due to the scarcity of datasets with high-quality joint annotations and mismatched control spaces across modalities. We present VidCRAFT3, a unified and flexible I2V framework that supports both independent and joint control over camera motion, object motion, and lighting direction by integrating three core components. Image2Cloud reconstructs a 3D point cloud from the reference image to enable precise camera motion control. ObjMotionNet encodes sparse object trajectories into multi-scale optical flow features to guide object motion. The Spatial Triple-Attention Transformer integrates lighting direction embeddings via parallel cross-attention. To address the scarcity of jointly annotated data, we curate the VideoLightingDirection (VLD) dataset of synthetic static-scene video clips with per-frame lighting-direction labels, and adopt a three-stage training strategy that enables robust learning without fully joint annotations. Extensive experiments show that VidCRAFT3 outperforms existing methods in control precision and visual coherence. Code and data will be released. Project page: https://sixiaozheng.github.io/VidCRAFT3/.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Image-to-video (I2V) generation is a powerful technique that brings still images to life, with broad applications across content creation, advertising, and animation. Controllable I2V generation, in particular, aims to animate a reference image according to user-specified control signals, such as text, object motion, and camera motion, while preserving high visual fidelity. Recent advancements in diffusion-based generative models and extensive web-scale data [1] have significantly enhanced the ability to generate temporally coherent and visually compelling videos from limited inputs, such as a single image or sparse annotations [2, 3, 4, 5, 6, 7]. However, achieving precise, simultaneous control of camera motion, object motion, and lighting direction remains a significant challenge in real-world scenarios.
@@
Existing approaches typically address these control signals independently or in a partially integrated manner. Methods focusing exclusively on camera control, such as CameraCtrl [7] and CamTrol [8], fail to adequately handle detailed object motion or dynamic lighting variations. Similarly, object-motion-centric frameworks like DragAnything [9], lack integrated capabilities for precise camera and lighting control. For lighting control, techniques like DILightNet [10] and NeuLighting [11] enable lighting adjustments. However, these methods often lack generalizability due to their reliance on specific categories, such as human faces, and the use of HDR maps. In addition, the dependence on control signals such as HDR maps [12] and background videos [13] restricts interactive control of lighting direction, making these methods less in open-domain scenarios.
@@
Simultaneous and precise control of camera motion, object motion, and lighting direction introduces multiple significant challenges: (1) Accurate camera motion control from a reference image requires reliable 3D priors to prevent drift and geometric distortion under large viewpoint changes. (2) Realistic and detailed object motion control demands effective representation of sparse object trajectories (e.g., a sequence of 2D keypoints) without compromising visual fidelity. (3) Dynamic lighting control necessitates integrating illumination adjustments coherently with both camera and object motion to ensure temporal consistency and visual realism. (4) High-quality datasets with joint annotations for camera motion, object motion, and lighting direction remain scarce.
@@
To overcome these challenges, we propose VidCRAFT3, which integrates three core components. First, the Image2Cloud module leverages DUSt3R [14] to reconstruct a 3D point cloud from a single reference image, enabling precise camera motion control by rendering the point cloud along user-defined camera trajectory. Second, ObjMotionNet encodes sparse object trajectories by extracting multi-scale motion features from Gaussian-smoothed optical flow maps to guide realistic object motion. Third, Spatial Triple-Attention Transformer integrates lighting embedding with image-text embeddings through parallel cross-attention layers. Capabilities of these modules are illustrated in Fig. 1, highlighting VidCRAFT3’s control over camera motion, object motion, and lighting direction. To address data scarcity, we introduce the VideoLightingDirection (VLD) dataset, providing synthetic static-scene video clips that are highly realistic, along with accurate per-frame lighting annotations. Additionally, we develop a three-stage training strategy that enables effective learning without fully joint annotations.
@@
The main contributions of this paper are: (1) To the best of our knowledge, VidCRAFT3 is the first I2V framework to achieve simultaneous control over camera motion, object motion, and lighting direction through a disentangled architecture combining Image2Cloud, ObjMotionNet, and Spatial Triple-Attention Transformer. (2) The VLD dataset provides synthetic videos with accurate per-frame lighting direction annotations, effectively addressing the scarcity of datasets with fully joint annotations. (3) A three-stage training strategy enables robust multi-element control without requiring dataset annotated with camera, object, and lighting signals. (4) Extensive experiments demonstrate that VidCRAFT3 achieves state-of-the-art performance surpassing existing methods in terms of control precision, visual quality, and generalization capabilities.
II Related Work
II-A Image-to-video Generation
@@
Image-to-video (I2V) generation [2, 3, 15, 4, 5, 16] aims to animate static images into dynamic videos while preserving visual content and introducing realistic motion. Recent advances in diffusion models [17, 18] have revolutionized video generation by extending pre-trained Text-to-Image (T2I) models like AnimateDiff [2] to incorporate temporal dimensions for motion generation. These methods integrate the input image as a condition, either through CLIP-based [19] image embeddings or by concatenating the image with noisy latent. For example, VideoCrafter1 [3], DynamiCrafter [4], and I2V-Adapter [5] use dual cross-attention layers to fuse image embeddings with noisy frames, ensuring spatial-aligned guidance. Similarly, Stable Video Diffusion (SVD) [20] replaces text embeddings with CLIP image embeddings, maintaining semantic consistency in an image-only manner. Another line of work, exemplified by SEINE [21], DynamiCrafter [4] and PixelDance [22], expands the input channels of diffusion models to concatenate the static image with noisy latents, effectively injecting image information into the model. However, these methods preserve input image fidelity during dynamic video generation but often struggle with fine-grained details due to reliance on global conditions.
II-B Motion-controlled Video Generation
Motion-controlled video generation [6, 23] focuses on creating high-fidelity videos with user-defined motion dynamics. Recent video generation models [24, 25, 26, 27, 28, 29] achieve impressive visual quality, but rely only on text and/or images for controlling content and motion, leading to coarse and implicit motion control. Existing approaches can be broadly divided into camera motion control, object motion control, and joint motion control. In the domain of camera motion control, one line of work conditions models on explicit camera intrinsics and extrinsics or on Plücker embeddings [7, 30, 31, 32, 33, 34, 35, 6, 36, 37]. For instance, MotionCtrl [6] injects extrinsic matrices into temporal attention to modulate viewpoint, whereas CameraCtrl [7] utilizes Plücker embeddings to incorporate geometric structure. Although these methods show promising results, directly mapping camera parameters to generative dynamics often restricts precision and generalization, particularly for trajectories outside the training distribution. To improve geometric consistency, another line of research lifts a reference image into 3D using depth maps or point clouds [38, 8, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Representative examples include ViewCrafter [44] and CamTrol [8], which render partial frames from reconstructed point clouds as guidance during generation. A further research direction explores training-free solutions [8, 48, 49, 50, 51] that derive motion representations by inverting temporal attention maps in pretrained models, as exemplified by MotionMaster [48] and MotionClone [50].
For object motion control, prior work explores complementary control signals and injection schemes. Flow-conditioned methods steer dynamics with sparse optical flow derived from user drags [52, 53, 54], as exemplified by DragNUWA [23], Image Conductor [55], and ReVideo [56]. Dense motion fields further refine pixelwise trajectories in MOFA-Video [57] and Motion-I2V [58]. Region-level cues provide a low-overhead interface, with bounding-box trajectories guiding objects[59, 60] in Boximator [61] and TrailBlazer [62], MagicMotion [63]. Region-level cues offer a low-overhead interface, where bounding-box trajectories guide object motion [59, 60, 62, 64], as demonstrated by Boximator [61] and MagicMotion [63]. To improve geometric fidelity, recent work leverages keypoint/point-map, mask-aware, and 3D cues. TrackGo [65] injects free-form masks and arrows via a lightweight temporal adapter. Tora [66] encodes trajectory maps with a 3D VAE. LeViTor [67] fuses depth with clustered keypoints for precise 3D control. Segmentation-aware schemes align user intent with object extent, with DragEntity [68] and DragAnything [9] mapping user drags to the corresponding masked entities. In parallel, training-free variants steer inference via attention guidance, energy-based denoising, or latent edits, enabling box/mask-conditioned control without finetuning [69, 70, 71].
@@ II-B Motion-controlled Video Generation
Current research on Joint Motion Control [72, 6, 73, 74, 75, 76, 77, 78], aiming to simultaneously control both camera and object motions, remains limited. Perception-as-Control [74] uses a simplified 3D scene to produce spatially aligned control signals; Motion Prompting [73] encodes spatiotemporal directives as structured prompts; MotionCanvas [77] converts 3D intent to 2D conditioning for diffusion models; CineMaster [78] combines 3D boxes and projected depth with a camera adapter; MotionAgent [76] decomposes text into camera extrinsics and object trajectories and composes them as optical flow. Unlike existing methods, we propose VidCRAFT3, the first framework to achieve simultaneous control over camera motion, object motion, and lighting direction. By combining 3D point cloud rendering, trajectory learning, and Spatial Triple-Attention Transformer, our approach effectively decouples these elements, ensuring temporal consistency and enhanced realism in complex scenes.
II-C Lighting-controlled Visual Generation
@@
Lighting-controlled visual generation aims to manipulate illumination while preserving scene geometry and materials. Previous methods primarily focus on portrait lighting [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89], laying the foundation for effective and accurate illumination modeling. Recent advances in diffusion models significantly improve the quality and flexibility of lighting control. Methods like DiLightNet [10] and GenLit [90] achieve fine-grained and realistic relighting through radiance hints and SVD, respectively. Facial relighting methods, including DifFRelight [87] and DiFaReli [88], produce high-quality portrait images. Frameworks like NeuLighting [11] focus on outdoor scenes using unconstrained photo collections, while GSR [91] combines diffusion models with neural radiance fields for realistic 3D-aware relighting. IC-Light [12] proposes imposing consistent light transport during training. Extending lighting control to video [89] introduces challenges such as temporal consistency and dynamic lighting effects. Recent techniques leverage 3D-aware generative models for temporally consistent relighting, as seen in EdgeRelight360 [92] and ReliTalk [93]. Neural rendering approaches [94, 95] use datasets like dynamic one-light-at-a-time (OLAT) for high-quality portrait video relighting, while reflectance field-based methods [96] infer lighting from exemplars. Light-A-Video [97] achieves training-free, temporally consistent video relighting via a Consistent Light Attention module and Progressive Light Fusion, while RelightVid [13] employs a 3D UNet with temporal layers trained on LightAtlas to deliver temporally consistent, high-quality results under flexible conditions (text, background videos, HDR maps) without intrinsic decomposition. Despite these advances, most prior work emphasizes portraits and HDR lighting conditions, limiting interactive control in general scenes. In contrast, VidCRAFT3 targets open, non-portrait scenarios and provides interactive, direction-level lighting control with geometry/material preservation and a scene-agnostic control interface.
III Method
III-A Overview
@@
Figure 2: Architecture of VidCRAFT3 for controlled image-to-video generation. The model builds on Video Diffusion Model (VDM) and consists of three main components: the Image2Cloud reconstructs 3D point cloud from a single reference image and generates point cloud renderings along a user-defined camera trajectory; the ObjMotionNet injects object dynamics into the UNet by encoding sparse trajectories into multi-scale motion features; the Spatial Triple-Attention Transformer integrates image, text, and lighting information via parallel cross-attention modules. The model enables I2V generation conditioned on arbitrary combinations of camera motion, object motion, and lighting direction.
VidCRAFT3 is, to the best of our knowledge, the first image-to-video (I2V) diffusion framework that enables both independent and joint control of camera motion, object motion, and lighting direction, as illustrated in Fig. 2. Given a reference image $I_{\text{ref}}\in\mathbb{R}^{H\times W\times 3}$ and a text prompt, the model generates a video $I=\{I_{f}\}_{f=1}^{F}$ with $I_{1}=I_{\text{ref}}$. The video frames will be generated according to the text description and all the control signals provided. Control signals can be used singly or in combination, namely camera trajectory $E=\{E_{f}\}_{f=1}^{F}$ with $E_{f}\in SE(3)$, sparse object trajectories $\mathcal{T}=\{\mathbf{s}_{n}^{f}=(x_{n}^{f},y_{n}^{f})\}$ for up to $N$ objects, and lighting direction $L_{\text{ref}}\in\mathbb{R}^{3}$.
The model builds on the I2V model DynamiCrafter [98] (Sec. III-C). Appearance is preserved by encoding $I_{\text{ref}}$ with the CLIP image encoder and injecting the embeddings through image cross-attention. Precise camera motion control is achieved by reconstructing a 3D point cloud from $I_{\text{ref}}$ with Image2Cloud and rendering along the camera trajectory $E$ to obtain geometry-aware renderings $R=\{R_{f}\}_{f=1}^{F}$. These renderings are encoded by VAE and concatenated with noise as the UNet input, which anchors global camera motion while allowing the video diffusion model (VDM) to refine appearance and temporal coherence. Object motion signal is provided by ObjMotionNet, which converts sparse trajectories into a dense smoothed motion tensor and injects multi-scale motion features into UNet encoder blocks so that structure is aligned without overconstraining details. Lighting direction is handled by a Spatial Triple Attention Transformer that integrates image, text, and lighting direction through parallel cross-attention layers. To address the scarcity of real-world datasets with fully joint annotations, we construct three specialized datasets (Sec. III-D) and adopt a three-stage training strategy (Sec. III-E) to progressively optimize the model.
III-B Preliminary
Video diffusion models (VDMs) represent a class of generative models that extend the principles of image diffusion to the domain of video generation. These models operate by defining a forward diffusion process that gradually transforms an initial video sample $x_{0}\sim p_{\text{data}}(x)$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ timesteps. The reverse process, parameterized by a denoising network $\epsilon_{\theta}(x_{t},t)$, learns to iteratively denoise the noisy latent representation $x_{t}$ to recover the original data $x_{0}$. The training objective is formulated as:
$\min_{\theta}\mathbb{E}_{t,x,\epsilon\sim\mathcal{N}(0,I)}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2}\right],$
To address the computational challenges associated with high-dimensional video data, Latent Diffusion Models (LDMs) are often employed. In this framework, a video $x\in\mathbb{R}^{F\times 3\times H\times W}$ is first encoded into a lower-dimensional latent space $z=\mathcal{E}(x)$, where $z\in\mathbb{R}^{F\times C\times h\times w}$. The diffusion and denoising processes are then performed in this latent space, significantly reducing computational complexity. The denoising process is conditioned on additional inputs $\mathbf{c}$, such as text prompts or motion control signals, enabling the generation of videos that adhere to specific semantic or temporal constraints. The final video is reconstructed through a decoder $\hat{x}=\mathcal{D}(z)$.
III-C Model Architecture